OpenStudy (anonymous):

Two opposite angles of a parallelogram are(5x-8) and (2x-82).find the measures of each angle of the parallelogram. Please help! urgent!!!!!

OpenStudy (anonymous):

opposite angles of a parallelogram are congruent. thus 5x-8=2x-82 x=-74/3

OpenStudy (anonymous):

Opposite angles of a parallelogram are congruent. ===> 5x-8 = 2x-82 ===> solving for x ===> 3x=-74 ==> x=-74/3

OpenStudy (anonymous):

===> the angle = 5*(-74/3)-8= -131.33

OpenStudy (anonymous):

how can angle be negative????

OpenStudy (anonymous):

but it seem that the negative sign is becasue of the given data

OpenStudy (anonymous):

maybe data is wrong

OpenStudy (anonymous):

so, it's quesionable here to check the given data

OpenStudy (anonymous):

I think he means sumthing other than opp angels

OpenStudy (anonymous):

that's right

OpenStudy (anonymous):

could be or some signs are revesed

OpenStudy (anonymous):

yaa jangzal tell us wat happened

OpenStudy (anonymous):

@Janzal are you around?.

OpenStudy (anonymous):

ya hey.. this was exactly how the question was worded..i think there might be some mistake in the question itself because this is an 8th grade syllabus (my brother's)

OpenStudy (anonymous):

yes, there some mistakes definitely